Solve by using the quadratic formula.
5x^2 + x = 3

Answer by malakumar_kos@yahoo.com(315)

Solve by using the quadratic formula.
5x^2 + x = 3
5x^2+x-3 = 0
compare with the std form we get a = 5,b = 1 and c = -3
use
x = -b+-sqrtb^2-4ac/2a
x = -1+-(1^2-4*5*-3)/2*5 = -1+-1+60/10 = -1+-sqrt 61/10

x = (-1+-sqrt 61)/10